OpenStudy (anonymous):

rationalize the denominator x2 square root 2x + 5

OpenStudy (anonymous):

\[\frac{ x ^{2} }{ \sqrt{2x+5} }\] so we need to multiply the fraction by 1 but one is square root 2x+5 therefore \[\frac{ x ^{2} }{ \sqrt{2x+5}} \times \frac{ \sqrt{2x+5} }{\sqrt{2x+5} }\]

OpenStudy (anonymous):

will you simplify that for me?